170525 UCSC

Serious marine drizzle event lasting into the afternoon.  Unfortunately, the RH sensor died (was getting dripped on by the tarp) and the Beefcake relay also died.  Both have been replaced.  11.8 mL of fog water collected with fan on manual.

170523 UCSC

Nice fog event last night.  The active collector opened at 6 am and about 60 mL of sample collected at 10:30 am. Collector cleaned with soap, water and dried, then blank taken at 2:30 pm.